OpenStudy (anonymous):

If f(x) = x+3/3 , then what is f-1(x)

OpenStudy (helder_edwin):

u mean \(f^{-1}(x)\) the inverse function?

OpenStudy (anonymous):

yes

OpenStudy (helder_edwin):

to find the inverse function u have to solve for \(x\) the following equation \[ \large y=f(x)=\frac{x+3}{3} \]

OpenStudy (helder_edwin):

try it!

OpenStudy (anonymous):

i still don't understand

OpenStudy (helder_edwin):

when u have y=f(x) u get a value for y every time you plug in some value for x.

OpenStudy (helder_edwin):

for instance if x=2 then \[ \large y=f(2)=\frac{2+3}{3}=\frac{5}{3} \]

OpenStudy (helder_edwin):

the idea is to the same the other way around. given a value of y find the corresponding (if it exists) value of x

OpenStudy (helder_edwin):

that's the inverse and to get that u have to solve the equation i told u

OpenStudy (anonymous):

oh okay thank you

OpenStudy (helder_edwin):

solve the equation and tell me what u get

OpenStudy (anonymous):

is it 1 over 3x plus 1?

OpenStudy (helder_edwin):

no

OpenStudy (helder_edwin):

u have the equation \[ \large y=\frac{x+3}{3} \] try to isolate \(x\)

OpenStudy (helder_edwin):